OpenStudy (anonymous):

As an incentive plan a company stated that employees who work for the company for 4 years would receive $516 and a laptop. Mr. Rodriquez worked for the company for 3.5 years. The company prorated the incentive plan and he still received the laptop, but only $264. What is the value of the laptop?

OpenStudy (triciaal):

think about equating with y = m x + b where the incentive received = y the slope m = amount in $ for each year worked and b would be the laptop

OpenStudy (triciaal):

x is the number of years worked

OpenStudy (anonymous):

ok so 516 = 4x + b 264 = 3.5x + b now should I eliminate the b's?

OpenStudy (anonymous):

I got $1500 is that right?

OpenStudy (triciaal):

2nd thought 516 = 4x + c 264 = 3.5x +c x = $/yr and c = laptop

OpenStudy (triciaal):

got the same

OpenStudy (anonymous):

thank you!

OpenStudy (triciaal):

you are welcome